OpenStudy (anonymous):

Write the decimal as a fraction or a mixed number in simplest form: 5.76

OpenStudy (campbell_st):

there are 2 decimal places so you will need to start by writing it over 100 so you get \[\frac{576}{100}\] which can be split into \[\frac{500}{100} + \frac{76}{100} ... or .....5 \frac{76}{100}\] now you need to simplify 76/100 by dividing both numerator and denominator by common factors. hope this helps

OpenStudy (anonymous):

so the answer would be 0.76 :/ im still confused

OpenStudy (campbell_st):

ummm I think you have a slight error @rachel1234 144 x 5 is not 576

OpenStudy (anonymous):

oops

OpenStudy (campbell_st):

ok... so the mixed number will be 5 76/100 look at 76/100 they are both even so you can divide the top and the bottom by 2 what would you get...?

OpenStudy (anonymous):

it's 144/25

OpenStudy (campbell_st):

the 5 will not change... its just the fraction part you need to simplify

OpenStudy (anonymous):

thank you i think i understand now